A 0.1 mol/L hydrochloric acid solution has a pH of 1.0 while a 0.1 mol/L hydrofluoric acid solution has a pH of 2.1. If these two acids have the same concentration, why is there a difference in the pH of these solutions?
Answer #1

HCl is a strong acid and it dissociates 100% completely in aqueous solutions therby produces 100% H+ ions, hence we can use formula pH = -log[H+],    pH = -log[0.1] = 1

whereas HF is a weak acid it will not completely dissociates into ions (not 100%) in aqueous solutions, hence we use Ka, dissociation constant which is different for every weak acid. for HF Ka = 7.2 X 10-4 now we have to use formula

[H+] = sqrt( Ka X C)

[H+] = sqrt ( 7.2 X 10-4X 0.1) = 8.5 X 10-3

pH = -log[H+] = -log[8.5 X 10-3] = 2.1